OpenStudy (anonymous):

how to find an equation in the form y-k=a(x-h)^2 if the x intercepts are -2 and 6 and the y-intercept is -6?

jimthompson5910 (jim_thompson5910):

if you know the two x-intercepts, then you can add them up and divide by 2 to get the value of h

OpenStudy (anonymous):

-2+6 = 4/2 = 2 h=2

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

the y intercept is -6 so x = 0 and y = -6 y-k=a(x-h)^2 y-k=a(x-2)^2 ... plug in h = 2 -6-k=a(0-2)^2 ... plug in (x,y) = (0,-6) -6-k = a(-2)^2 -6 - k = 4a -6 - k+6 = 4a+6 -k = 4a + 6 k = -4a - 6 Now let's plug in an x intercept, say (x,y) = (-2,0) y-k=a(x-2)^2 0-k=a(-2-2)^2 -k = a(-4)^2 -k = 16a k = -16a -4a - 6 = -16a ... replace k with -4a-6 now solve for 'a' to get a = ???

OpenStudy (anonymous):

a=1/2

jimthompson5910 (jim_thompson5910):

yes, a = 1/2 k = -4a - 6 k = -4(1/2) - 6 k = ???

OpenStudy (anonymous):

k=-8 is the final equation y+8=(1/2)(x-2)^2?

jimthompson5910 (jim_thompson5910):

yes it is

OpenStudy (anonymous):

Oh, thank you very much!

jimthompson5910 (jim_thompson5910):

you're welcome